# The Relative Contributions of Accretion Disk versus Jet to the Optical and Mid-infrared Variability of Seyfert Galaxies

### v1

@@ 4 Results and discussion
On the other hand, as discussed in Sect. 3.4, studying the color variability of AGNs can provide an effective way to disentangle between quasi-thermal emission from the accretion disc and non-thermal synchrotron emission from the jet, which contributes to the overall observed emission. Therefore, for our samples of gNLS1s, ngNLS1s and gBLS1s, we generated color-magnitude (g-r vs. r & r-i vs. i) diagrams in optical regime and W1-W2 vs. W2 in MIR regime using their quasi-simultaneous data observed in respective bands. In Fig. 5, long-term color-magnitude plots for (g-r) vs. r, (r-i) vs. i, & (W1-W2) vs. W2, respectively, are presented in top, middle, and bottom for representative examples of the gNLS1-J130522.74511640.2 (left), ngNLS1-J144019.82502201.7 (middle), and gBLS1-J110249.85525012.7 (right) from the current sample, illustrating RWB trend both in optical and MIR wavelengths, respectively. In these figures, the black dashed line represents the best fit, obtained using orthogonal distance regression, accounting for uncertainties in both color and magnitude measurements.
Furthermore, considering the criterion of and set to identify stronger BWB and RWB, respectively from the color-magnitude diagrams, we have found 35% (6/17) and 61% (78/128) of gNLS1s and ngNLS1s, respectively, exhibiting RWB trend in (g-r) vs. r color-magnitude diagram (see Table 3), however, only a gNLS1-J150506.48032630 6% (1/17) showed BWB (see Table 4). Such an RWB trend is found for 50% (5/10) and 74% (55/74) in samples of gNLS1s and ngNLS1s, respectively, when considering the color-magnitude diagram (r-i) vs. i (see Table 3). Similarly, in the sample of gBLS1s, 80% (8/10) and 50% (3/6) exhibit the RWB trend in the color-magnitude diagrams (g-r) vs. r and (r-i) vs. i, respectively (see Table 3). Note that 10 out of 17 gNLS1s, 50 out of 128 ngNLS1s, and 2 out of 10 gBLS1s didn’t show strong correlation between (g-r) color and r-band magnitude (see Table 3). However, for the correlation between the (r-i) color and i-band magnitude, these numbers are 5 out of 10 gNLS1s, 19 out of 74 ngNLS1s, and 3 out of 6 gBLS1s (see Table 3).
@@
On the other hand, similar to the optical wavelength, 68% (15/22), 96% (169/176), and 30% (3/10) of gNLS1s, ngNLS1s, and gBLS1s, respectively, showed the RWB trend in color magnitude diagram of (W1-W2) vs. W2 (see Table 3). Contrary to the RWB trend exhibited by most sources, only a gBLS1s-J122454.45212246.3 showed a BWB trend (see Table 3). The remaining 32% (7/22), 4% (7/176), and 60% (6/10) of gNLS1s, ngNLS1s, and gBLS1s, respectively, didn’t exhibit a strong correlation between (W1-W2) color and W2-band magnitude (see Table 3).
@@
The so-called BWB and RWB color trends from the AGN in the optical wavelength attribute to domination of non-thermal emission from the jet over thermal emission from the accretion disc, and vice-versa, respectively (e.g., see Malkan, 1983; Pian et al., 1999; Rani et al., 2010; Sakata et al., 2010; Ikejiri et al., 2011; Mao & Zhang, 2016; Negi et al., 2022). The BWB trend that predominantly appeared in the BL Lacs class of AGN can be explained with the shock-in-jet model (see Türler, 2011) where electrons are accelerated to higher energies at the front of the shock, and since high-energy electrons lose their energy faster compared to low-energy electrons by radiative cooling process, thus making the high-energy band (in the current work g-band compared to r-band and r-band compared to i-band) more variable (Kirk et al., 1998; Mastichiadis & Kirk, 2002). It could also be explained by an energy injection of fresh electrons due to internal shocks in the relativistic shell into the emitting regions, leading to an increase in the number of high energy electrons and eventually shifting to higher energy (e.g., Spada et al., 2001; Zhang et al., 2002; Fiorucci et al., 2004). Apart from the energy injection scenario, the BWB trend can also be attributed to variations in the beaming factor, , of the emitting region. Changes in not only result in an apparent increase in observed flux, following the relation , but also cause a shift in the observed frequency, (e.g., Villata et al., 2004; Papadakis et al., 2007; Larionov et al., 2010).
@@
On the other hand, RWB trend that appeared in the Seyfert galaxies, can be explained by the domination of the disc’s fluctuations over the jet’s fluctuations. Since long-term optical variability in Seyfert galaxies is exhibited by a mix of disc and jet emissions, if the accretion disc’s fluctuations are primarily responsible for the variability but jet contributes more in blue/UV, then any brightening of the disc may increase the domination of redder thermal emission, leading to RWB trend (e.g., see Malkan, 1983; Pian et al., 1999; Sakata et al., 2010). The RWB trend observed in Seyfert galaxies and FSRQs can also be accommodated in the shock-in-jet model. When a shock propagates through a jet and interacts with regions of higher electron density, stronger magnetic fields, or turbulence, the emission from that region intensifies. Initially, higher-energy radiation emerges first, followed by a delayed peak at lower frequencies (e.g. Valtaoja et al., 2002). In the later stages of a flare, the flux enhancement in the redder bands may become more pronounced as the contribution from shorter wavelengths stabilizes or declines, leading to an overall RWB trend. This effect is particularly relevant in Seyfert galaxies, where the jet emission is weaker, and contributions from the disc and dust are more significant. In FSRQs, the strong thermal emission from the accretion disc, which peaks in UV, plays a crucial role. As the synchrotron emission from the jet increases, typically peaks at lower frequencies, it can shift the overall spectral energy distribution to redder colors, reinforcing the RWB trend (e.g., see Gu et al., 2006).
@@
Thus, based on the long-term optical color variability observed in the current samples of gNLS1s, ngNLS1s, and gBLS1s, the dominant mechanism driving their long-term variability appears to be associated with processes within the accretion disc. This variability is likely induced by instabilities within the accretion disc. However, on shorter timescales, particularly on the intra-day timescale, the observed variability may be primarily attributed to jet-related processes.
@@
On the other hand, based on the mid-infrared color-magnitude analysis (W1-W2) vs. W2, a large percentage 68%, 96%, and 30%, of gNLS1s, ngNLS1s, and gBLS1s, respectively, exhibit an RWB trend similar to that exhibited in optical wavelength (see Table 3). The observed RWB trend in the MIR wavelength in the current sample of Seyfert galaxies can be attributed to the dominance of a variable nuclear component over the constant host galaxy component, where the increased brightness of the AGN leads to a stronger contribution from the AGN dust torus to the observed MIR emission, resulting in a redder (W1-W2) color (e.g., see Stern et al., 2012; Yang et al., 2018). Additionally, a similar color trend observed in optical and mid-infrared wavelengths for most of the sources also reconfirms the reprocessing scenario. Thus, long-term MIR color variability analysis of gNLS1s, ngNLS1s, and gBLS1s reconfirms the conclusions drawn based on optical wavelength for the current sample.
@@ 5 Conclusions
(i) The highest variability is exhibit by gBLS1s in both optical (g, r, i bands) and mid-infrared (W1, W2 bands), followed by gNLS1s, while ngNLS1s show the least variability across both wavelengths. The higher optical and mid-infrared variability observed in gBLS1s compared to gNLS1s suggests a more closely aligned jet toward the observer line of sight in gBLS1s. This could also be due to comparatively lower accretion rates of gBLS1, which contaminate less effectively non-thermal Doppler-boosted synchrotron emission of gBLS1s as compared to gNLS1s. However, the least variability (DC of 5%) shown by ngNLS1s among different samples of Seyfert galaxies studied here suggests their long-term variability due to instabilities in the accretion
@@
(iii) A similar kind of variability nature resulted from the samples of gNLS1s, ngNLS1s, and gBLS1s in the mid-infrared W1 and W2 bands, supporting a reprocessing of optical/UV emission from the central engine of Seyfert galaxies by the dusty torus to the MIR wavelength. The reprocessing scenario was confirmed based on a significant correlation found between optical and mid-infrared variability amplitudes for gNLS1s and gBLS1s in the g and r bands, while this correlation is weaker in the i band and nearly absent in ngNLS1s, except for their i band data, which are very scarcely sampled. Thus, no hints of a reprocessing scenario have been found for ngNLS1s based on flux variability analysis, which could be due to a large contamination in the variability amplitude by the host galaxies.
(iv) Long-term color-magnitude analysis reveals that most gNLS1s, ngNLS1s, and gBLS1s exhibit an RWB trend in both optical and MIR wavelengths. Hence, it suggests that the dominant mechanism driving their long-term variability appears to be associated with processes within the accretion disc. This variability is likely induced by instabilities within the accretion disc. However, on shorter timescales, particularly on the intra-day timescale, the observed variability may be primarily attributed to jet-related processes. However, we could not find any hint of a reprocessing scenario for ngNLS1s based on the nearly absent correlation found between optical and mid-infrared variability amplitudes, but a similar RWB color trend found for most of the sources in both optical and mid-infrared wavelengths strongly supports the reprocessing scenario.
(v) Thus, long-term mid-infrared flux and color variability analysis of gNLS1s, ngNLS1s, and gBLS1s reconfirms the conclusion drawn based on the optical wavelength for the current sample.
(vi) The present analysis of gNLS1s, ngNLS1s, and gBLS1s highlights the significance of multi-wavelength studies in disentangling the contributions of various emission mechanisms in Seyfert galaxies. The findings underscore the necessity of a more systematic approach involving larger sample sizes and high-cadence, multi-band observations across different timescales to refine our understanding of the physical processes governing AGN variability in future studies.

### v2

@@ 4 Results and discussion
On the other hand, as discussed in Sect. 3.4, studying the color variability of AGNs can provide an effective way to disentangle between quasi-thermal emission from the accretion disc and nonthermal synchrotron emission from the jet, which contributes to the overall observed emission. Therefore, for our samples of gNLS1s, ngNLS1s, and gBLS1s, we generated color-magnitude (g-r vs. r & r-i vs. i) diagrams in the optical regime and W1-W2 vs. W2 in the MIR regime using their quasi-simultaneous data observed in respective bands. However, during the color-magnitude analysis, we noticed that its trend reverses when we use shorter wavelength (bluer) magnitudes in place of longer wavelength (redder) on the X-axis, especially for ngNLS1s in the MIR wavelength. To account for this, we produced three combinations of color-magnitude diagrams using m1 and m2 magnitudes, that are, (m1-m2) vs. m2, (m1-m2) vs. m1, and (m1-m2) vs. (m1+m2). In Fig. 6, Top: long-term (g-r) color variation versus r-magnitude, g-magnitude and (g+r) (magnitude) plots for representative example of the gNLS1-J094857.32002225.5 from the current sample. Middle: Same as top but for the (r-i) color vs. i, r and (r+i) magnitudes for the same target. Bottom: Same as top, but for the (W1-W2) color vs. W2, W1 and (W1+W2) magnitudes for the same target. All nine panels illustrate negative trends in the color-magnitude diagrams in both optical and mid-infrared wavelengths. The black dashed line represents the best-fit curve obtained using ODR, which accounts for uncertainties in both color and magnitude measurements. The best-fit linear regression parameters, including the slope, intercept, and Pearson’s correlation coefficient (Pearson r), are displayed in the upper left corner of each panel. A very similar color-magnitude diagrams are presented for ngNLS1-J144153.98482227.2 and gBLS1-J110249.85525012.7 in the Figs. 7 & 8, respectively.
Since a genuine correlation should not depend on the specific choice of magnitude in a color–magnitude diagram, the observed reversal of trends in cases where the disk contribution dominates led us to base our conclusions primarily on the (m1-m2) color versus (m1+m2) magnitude diagrams. We further noted that, for this method, the strength of the correlation is consistently weaker compared to the other two combinations, even in cases of true correlations (i.e., those without trend reversals). Therefore, we adopt a more relaxed threshold, requiring , to define strong and reliable correlations. Under this criterion, corresponds to a BWB trend, while indicates a RWB trend.
Applying this framework to the (g-r) vs. (g+r) diagrams, we find that (4/17) and (5/17) of gNLS1s exhibit RWB and BWB behavior, respectively (see Table 3). For ngNLS1s, the corresponding fractions are (13/128) and (31/128), respectively. In the (r-i) vs. (r+i) diagrams, we find (5/10) and (1/10) of gNLS1s exhibiting RWB and BWB behavior, respectively, while the ngNLS1 sample yields (36/74) RWB and (13/74) BWB trends. For gBLS1s, (8/10) exhibit the RWB trend in (g-r) vs. (g+r), and (3/6) exhibit the same trend in (r-i) vs. (r+i), with the exception of J094420.44613550.1, which shows a BWB trend in the latter diagram. We also note that a significant fraction of sources show no strong correlations under the adopted criterion. Specifically, 8 out of 17 gNLS1s (), 84 out of 128 ngNLS1s (), and 2 out of 10 gBLS1s () do not exhibit strong correlations between (g-r) color and (g+r) magnitude. For the (r-i) vs. (r+i) diagrams, the corresponding fractions are 4 out of 10 gNLS1s (), 25 out of 74 ngNLS1s (), and 2 out of 6 gBLS1s () (see Table 3).
@@
On the other hand, similar to the optical wavelength, 55% (15/22), 28% (49/176), and 30% (3/10) of gNLS1s, ngNLS1s, and gBLS1s, respectively, showed the RWB trend in color magnitude diagram of (W1-W2) vs. (W1+W2) (see Table 3). Contrary to the RWB trend exhibited by most sources, a smaller percentage of 4% (1/22), 4% (7/176), and 30% (3/10) of gNLS1s, ngNLS1s, and gBLS1s, respectively, showed the BWB trend (see Table 3). The remaining 41% (9/22), 67% (118/176), and 40% (4/10) of gNLS1s, ngNLS1s, and gBLS1s, respectively, didn’t exhibit a strong correlation between (W1-W2) color and (W1+W2) magnitude (see Table 3).
Based on the color–magnitude analysis (m1-m2) vs. (m1+m2) for gNLS1s, ngNLS1s, and gBLS1s across optical to MIR wavelengths, we find the following trends: in the shorter wavelength regime, BWB behavior dominates over RWB in the color–magnitude diagrams (e.g., (g-r) vs. (g+r)). In contrast, in the longer wavelength regimes, RWB behavior dominates over BWB (e.g., (r-i) vs. (r+i) and (W1-W2) vs. (W1+W2)), except for gBLS1s, where RWB behavior is dominant from the optical through the MIR.
@@
The so-called BWB and RWB color trends from the AGN in the optical wavelength attribute to domination of nonthermal emission from the jet over thermal emission from the accretion disc, and vice-versa, respectively (e.g., see Malkan, 1983; Pian et al., 1999; Rani et al., 2010; Sakata et al., 2010; Ikejiri et al., 2011; Mao & Zhang, 2016; Negi et al., 2022). The BWB trend that predominantly appeared in the BL Lacs class of AGN can be explained with the shock-in-jet model (see Türler, 2011) where electrons are accelerated to higher energies at the front of the shock, and since high-energy electrons lose their energy faster compared to low-energy electrons by radiative cooling process, thus making the high-energy band (in the current work g-band compared to r-band and r-band compared to i-band) more variable (Kirk et al., 1998; Mastichiadis & Kirk, 2002). It could also be explained by an energy injection of fresh electrons due to internal shocks in the relativistic shell into the emitting regions, leading to an increase in the number of high energy electrons and eventually shifting to higher energy (e.g., Spada et al., 2001; Zhang et al., 2002; Fiorucci et al., 2004). Apart from the energy injection scenario, the BWB trend can also be attributed to variations in the beaming factor, , of the emitting region. Changes in not only result in an apparent increase in observed flux, following the relation , but also cause a shift in the observed frequency, (e.g., Villata et al., 2004; Papadakis et al., 2007; Larionov et al., 2010).
@@
On the other hand, the RWB trend that appeared in the Seyfert galaxies in the current work, especially towards longer wavelengths, can be explained by the domination of the disc’s fluctuations over the jet’s fluctuations. Since long-term optical variability in Seyfert galaxies is exhibited by a mix of disc and jet emissions, if the accretion disc’s fluctuations are primarily responsible for the variability but jet contributes more in blue/UV, then any brightening of the disc may increase the domination of redder thermal emission, leading to RWB trend (e.g., see Malkan, 1983; Pian et al., 1999; Sakata et al., 2010). The RWB trend observed in Seyfert galaxies and FSRQs can also be accommodated in the shock-in-jet model. When a shock propagates through a jet and interacts with regions of higher electron density, stronger magnetic fields, or turbulence, the emission from that region intensifies. Initially, higher-energy radiation emerges first, followed by a delayed peak at lower frequencies (e.g. Valtaoja et al., 2002). In the later stages of a flare, the flux enhancement in the redder bands may become more pronounced as the contribution from shorter wavelengths stabilizes or declines, leading to an overall RWB trend. This effect is particularly relevant in Seyfert galaxies, where the jet emission is weaker, and contributions from the disc and dust are more significant. In FSRQs, the strong thermal emission from the accretion disc, which peaks in UV, plays a crucial role. As the synchrotron emission from the jet increases, typically peaks at lower frequencies, it can shift the overall spectral energy distribution to redder colors, reinforcing the RWB trend (e.g., see Gu et al., 2006).
@@
Thus, based on the long-term optical color variability observed in the current samples of gNLS1s, ngNLS1s, and gBLS1s, the dominant mechanism driving their long-term variability appears to be associated with processes within the accretion disc, which appears to be more prominent towards longer optical wavelengths. This variability is likely induced by instabilities within the accretion disc. However, on shorter timescales, particularly on the intra-day timescale, the observed variability may be primarily attributed to jet-related processes.
@@
On the other hand, based on the mid-infrared color-magnitude analysis (W1-W2) vs. W1+W2, a large percentage 55%, 28%, and 30%, of gNLS1s, ngNLS1s, and gBLS1s, respectively, exhibit an RWB trend similar to that exhibited in the optical color-magnitude diagram (r-i) vs. r+i (see Table 3). The observed RWB trend in the MIR wavelength in the current sample of Seyfert galaxies can be attributed to the dominance of a variable nuclear component over the constant host galaxy component, where the increased brightness of the AGN leads to a stronger contribution from the AGN dust torus to the observed MIR emission, resulting in a redder (W1-W2) color (e.g., see Stern et al., 2012; Yang et al., 2018). Additionally, a similar color trend observed in optical and mid-infrared wavelengths for most of the sources also reconfirms the reprocessing scenario.
@@ 5 Conclusions
(i) The highest variability is exhibit by gBLS1s in both optical (g, r, i bands) and mid-infrared (W1, W2 bands), followed by gNLS1s, while ngNLS1s show the least variability across both wavelengths. The higher optical and mid-infrared variability observed in gBLS1s compared to gNLS1s suggests a more closely aligned jet toward the observer line of sight in gBLS1s. This could also be due to the comparatively lower accretion rates of gBLS1, which less effectively contaminate the nonthermal Doppler-boosted synchrotron emission of gBLS1s compared to gNLS1s. However, the least variability (DC of 5%) shown by ngNLS1s among different samples of Seyfert galaxies studied here suggests their long-term variability due to instabilities in the accretion
@@
(iii) A similar kind of variability nature resulted from the samples of gNLS1s, ngNLS1s, and gBLS1s in the mid-infrared W1 and W2 bands, supporting a reprocessing of optical/UV emission from the central engine of Seyfert galaxies by the dusty torus to the MIR wavelength. The reprocessing scenario was confirmed based on a significant correlation found between optical and mid-infrared variability amplitudes for gNLS1s and gBLS1s in the g and r bands, while this correlation is found to be generally weak or statistically insignificant in the i band and nearly absent in ngNLS1s, except for a notable trend in the i band data, which are very scarcely sampled. Therefore, nearly statistically insignificant correlations are found for ngNLS1s, probably due to a combination of a small sample size and significant contamination of host galaxy light in the optical bands, rather than the absence of an intrinsic physical connection between optical and MIR variability. Furthermore, a strong correlation between optical and MIR variability amplitudes across different time scales supports the reprocessing scenario.
(iv) Long-term color-magnitude analysis reveals that most gNLS1s, ngNLS1s, and gBLS1s exhibit an RWB trend in both optical (towards longer wavelength side) and MIR wavelengths, strengthening the reprocessing scenario. Hence, based on flux and color variability analysis, it suggests that the dominant mechanism driving their long-term variability appears to be associated with processes within the accretion disc. This variability is likely induced by instabilities within the accretion disc. However, on shorter timescales, particularly on the intra-day timescale, the observed variability may be primarily attributed to jet-related processes. A similar RWB color trend found for most sources in both optical and mid-infrared wavelengths strongly supports the reprocessing scenario.
(v) The present analysis of gNLS1s, ngNLS1s, and gBLS1s highlights the significance of multi-wavelength studies in disentangling the contributions of various emission mechanisms in Seyfert galaxies. The findings underscore the necessity of a more systematic approach involving larger sample sizes and high-cadence, multi-band observations across different timescales to refine our understanding of the physical processes governing AGN variability in future studies.